OpenStudy (anonymous):

cubed square root of 7 multiplied by cubed square root of 9

OpenStudy (anonymous):

do you mean a cube root?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt[3]{7}*\sqrt[3]{9}= \sqrt[3]{7*9}\]

OpenStudy (anonymous):

is this considered a single radical as an answer?

OpenStudy (anonymous):

no cobe root of 63 is the radical

OpenStudy (anonymous):

cube*

OpenStudy (anonymous):

ok that is what i put as answer. thanks

OpenStudy (anonymous):

radical is the symbol \[\sqrt{}\]

OpenStudy (anonymous):

\[ \sqrt[4]{3}\times \sqrt[3]{2?}\]

OpenStudy (anonymous):

change them to the same exponent.